# Compositional Oil Spill Detection Based on Object Detector and Adapted Segment Anything Model from SAR Images

### v1

@@ III-D Qualitative Analysis
Fig. 4 demonstrates the YOLOv8-SAM and SAM-OIL perform well in scenarios where there are elongated dark areas and small dark areas with simple surroundings, such as those without look-alikes (e.g., first column of Fig. 4). Benefiting from the integration of OMF and Adapter, the segmentation of objects with fuzzy boundaries using the SAM-OIL is more aligned with the ground truth, as illustrated in the second and third columns. Nevertheless, some small-sized look-alikes tend to be misclassified as oil spills by the SAM-OIL, highlighting the difficulty of determining oil spills based solely on SAR images.
@@ IV Conclusion
We propose a composite framework, SAM-OIL, which integrates YOLOv8 with the SAM, incorporating OMF and Adapter modules. To our best knowledge, this is the first time the SAM has been applied to oil spill detection. Compared to existing oil spill detection methods, SAM-OIL achieves the highest mIoU, demonstrating the significant potential of SAM-based composite framework in oil spill detection. Furthermore, the version of SAM-OIL without the Adapter module, namely YOLOv8-SAM+OMF, achieves accuracy comparable to existing oil spill detection methods. It indicates that the SAM-OIL framework can reduce the need for finely annotated samples. Our experimental results also suggest that the object detector has a large contribution on the final accuracy. Therefore, prioritizing the development of object detectors could further improve the detection performance in future.

### v2

@@ III-D Qualitative Analysis
Fig. 4 shows that YOLOv8-SAM and SAM-OIL perform well in detecting elongated and small dark areas with simple surroundings, as seen in column 1. The integration of OMF and Adapter helps SAM-OIL segment objects with fuzzy boundaries more accurately, as shown in columns 2 and 3. However, small-sized look-alikes, particularly in the top left of column 3 and the bottom left of column 4, are sometimes misclassified as oil spills by the SAM-OIL, highlighting the challenges of using SAR imagery alone for oil spill detection.
@@ IV Conclusion
We proposed a composite framework, SAM-OIL, which integrates YOLOv8 with the SAM, incorporating OMF and Adapter modules. To the best of our knowledge, this is the first time the SAM has been applied in oil spill detection. Compared with existing methods, SAM-OIL achieves the highest mIoU on the M4D dataset with 69.52%, demonstrating the significant potential of the SAM-based composite framework in oil spill detection. Furthermore, the SAM-OIL version without the adapter module, YOLOv8-SAM+OMF, achieved 65.82% mIoU, indicating the framework’s ability on reducing the dependence on finely annotated samples.
Qualitative analysis shows that small-sized look-alikes are often confused with oil spills, suggesting the limitations in relying solely on SAR satellite imagery for detection and the need of ancillary data (e.g., oil platform locations, ship traffic)[3]. Within the SAM-OIL framework, the object detector significantly contributes to the accuracy improvement. Hence, prioritizing object detectors could further improve detection performance in the future.